OpenStudy (anonymous):

A population of bacteria is growing according to the function below, where t is in hours. How many hours will it take for the population to grow to 12,000 bacteria? Round your answer to the nearest integer. Do not include units in your answer. B(t) = 50 * e^2t

OpenStudy (anonymous):

solve \[12000=50e^{2t}\] for t 1) divide by 50 get \[240=e^{2t}\] 2)rewrite in equivalent logarithmic form as \[\ln(240)=2t\] 3)divide by 2 get \[\frac{\ln(240)}{2}\] then a calculator